A deck of cards contains 24 cards where 4 are red. If 6 cards are drawn at random without replacement from this deck, what is the probability that no red cards are drawn?


1
Expert's answer
2021-05-07T09:04:49-0400

6 cards out of 24 can be chosen in "n = C_{24}^6" ways.

6 cards out of 20 non-red can be chosen in "m = C_{20}^6" ways.

then wanted probability is "p = \frac{m}{n} = \frac{{C_{20}^6}}{{C_{24}^6}} = \frac{{20!}}{{6! \cdot 14!}} \cdot \frac{{6! \cdot 18!}}{{24!}} = \frac{{510}}{{1771}}"

Answer: "p = \frac{{510}}{{1771}}"